The radius of the planet Venus is nearly the same as that of earth, but it's mass is only 80% that of the earth.If an object weighs We on the earth , what does it weighs on Venus? Calculate the value of g on Venus


1
Expert's answer
2021-03-23T11:15:43-0400

a) We can find the weight of the object on the Earth as follows:


"W_E=mg_E=\dfrac{GM_Em}{R^2}."

Then, the weight of the object on the Venus will be:


"W_V=mg_V=\dfrac{GM_Vm}{R^2}=\dfrac{G0.8M_Em}{R^2}=0.8W_E."


b) Let's find the acceleration due to gravity on Venus:


"g_V=\dfrac{GM_V}{R_V^2}=\dfrac{G0.8M_E}{R_E^2}=0.8g_E=0.8\cdot9.8\ \dfrac{m}{s^2}=7.84\ \dfrac{m}{s^2}."